# Self-Filtered Distillation with LLMs-generated Trust Indicators for Reliable Patent Classification

### v1

@@ Abstract
Large language models (LLMs) increasingly generate natural language rationales to enhance interpretability, but these often contain logical errors, label mismatches, and domain-specific misalignments. Directly using such rationales as supervision risks propagating noise and undermining training stability. To address this challenge, we introduce Self-Filtered Distillation, a framework specifically tailored for patent classification, which treats LLM-generated rationales as trust signals rather than ground-truth supervision. The framework employs selective distillation guided by three unsupervised trust metrics: (1) Self-Consistency, which measures the stability of LLM-generated rationales across multiple generations; (2) Class Entailment Alignment, which assesses semantic coherence with patent-specific class definitions; and (3) LLM Agreement Scoring, which validates rationale–label plausibility. These metrics are integrated into a unified trust score that primarily weights training samples while optionally filtering out extremely low-trust cases, enabling reasoning-aware supervision. Experiments on the USPTO-2M dataset, a widely used benchmark for patent classification, show that our method outperforms label-based learning and conventional distillation in accuracy, stability, and interpretability, establishing a reliable paradigm for leveraging reasoning-aware trust indicators in patent analytics.
@@ 1 Introduction
Patent classification is critical for organizing large-scale patent corpora into technical classification schemes, serving as the foundation for applications such as intellectual property management, technology scouting, and innovation analysis (Jiang and Goetz, 2024; Yoo et al., 2025). Unlike general-domain texts, patents exhibit a multifaceted nature: highly specialized terminology, complex claim structures, and strategically ambiguous expressions, which pose significant challenges for automated classification. Our approach is specifically tailored to the patent domain, distinguishing it from general text classification techniques. Recent advances in large language models (LLMs) offer strong potential to address these challenges by processing massive patent datasets with improved accuracy and generating natural language rationales that enhance interpretability and provide a reasoning-based foundation for trust in classification outcomes (Lee and Hsiang, 2019; Bai et al., 2024).
Figure 1: Direct distillation vs. self-filtered distillation: Weighting and filtering rationales through trust metrics yields more stable and interpretable student models.
@@
Beyond predictive accuracy, research in NLP has increasingly emphasized explanation-driven learning as a means of enhancing interpretability and trustworthiness. This line of work leverages human-annotated rationales as auxiliary supervision signals, demonstrating that explanation can improve both generalization and interpretability (Camburu et al., 2018; DeYoung et al., 2020). While promising, such approaches incur high annotation costs and are often limited to short texts and simple classification tasks. Consequently, they fall short when applied to the classification of multifaceted and lengthy patents, where explanation-driven supervision faces domain-specific challenges.
@@
With the advent of LLMs, research has increasingly shifted toward leveraging model-generated rationales to strengthen learning without requiring expensive human annotation (Wei et al., 2022; Hsieh et al., 2023). While this paradigm offers scalability and interpretability, LLM-generated rationales often contain logical inconsistencies, label mismatches, and semantic misalignment with domain-specific classification schemes (Turpin et al., 2023), unsuitable for multifaceted patents. For instance, an LLM may justify categorizing a digital communication patent (H04L) based on antenna structures, or misclassify an artificial intelligence patent (G06N) as general data processing (G06F). Such errors extend beyond performance degradation: in the patent domain, they can lead to misguided technology mapping, flawed market analysis, or even intellectual property disputes. Treating all rationales as direct supervisory signals risks propagating these errors and reducing training stability. Therefore, high-stakes applications of patent classification demand a framework that can quantitatively assess rationale quality and selectively incorporate only trustworthy rationales (Lampinen et al., 2022).
@@
Empirical evaluations on the USPTO-2M dataset demonstrate that SFD surpasses traditional label-based learning and conventional distillation methods, achieving superior performance in predictive accuracy, training stability, and interpretability. This indicates that trust-based LLM rationale modulation not only enhances classification performance but also establishes a new learning paradigm for effectively leveraging LLM-generated rationales.
@@
  • Trust-based distillation paradigm: which treats LLM-generated rationales as trust indicators rather than ground-truth signals, reducing instability and error propagation while turning model explanations into reliable learning resources.
@@
  • Patent-domain validation: Applied to patent classification, the SFD outperforms label-based learning and conventional distillation in accuracy, stability, and interpretability, demonstrating its value for trustworthy patent analytics.
@@ 2.1 Patent Classification
Patent classification is a central task in intellectual property management, supporting prior art search, and legal decision-making. Early studies used TF–IDF features with classifiers such as SVMs or logistic regression, with limited performance due to domain-specific terminology and high-dimensional feature spaces (Fall et al., 2003).
Deep learning brings about significant improvements: CNN- and RNN-based models capture semantic and structural patterns from claims and abstracts, outperforming methods (Li et al., 2018). Still, the challenges of long documents, technical vocabulary, and concept interdependencies make patent classification harder than general text classification. Transformer-based models open a new paradigm. Domain-adapted encoders such as PatentBERT (Lee and Hsiang, 2019) and PatentSBERTa (Bekamiri et al., 2021) outperform BERTs, while generative approaches like PatentGPT show the potential of large-scale pretraining (Bai et al., 2024). Leveraging the hierarchical structure of the Cooperative Patent Classification (CPC) also helps address the complexity of multi-label classification (Zhou et al., 2021). Large-scale datasets like USPTO-2M, with millions of patents and thousands of CPC classes, further accelerate progress but also introduce issues of computational cost and severe label imbalance (Risch and Krestel, 2020).
@@ 2.2 Explanation-driven Learning
Chain-of-Thought (CoT) prompting enables LLMs to solve complex reasoning tasks by generating intermediate rationales, improving predictive performance and interpretability across arithmetic reasoning, commonsense inference, and logical deduction (Wei et al., 2022; Kojima et al., 2022; Wang et al., 2023). This has inspired the distillation research that reasoning traces from large teacher models are transferred into smaller students (Hsieh et al., 2023). For instance, KeyPoint-CoT compresses lengthy rationales into concise key points for efficiency and faithfulness (Feng et al., 2024). Yet most distillation approaches assume uniform rationale quality, which risks propagating inconsistent or erroneous explanations. This is especially problematic in patent classification, where misaligned rationales can distort label assignments.
Explanation-driven learning more broadly uses rationales as auxiliary supervision. Datasets like e-SNLI (Camburu et al., 2018) and ERASER (DeYoung et al., 2020) show that human-written explanations improve generalization and transparency, while in-context rationales aid robustness in few-shot learning (Lampinen et al., 2022; Wiegreffe et al., 2021). However, LLM-generated rationales often suffer from incompleteness, unfaithfulness, or contradictions with labels (Turpin et al., 2023). These issues are amplified in patent classification due to multifaceted patent features including lengthy claims, technical vocabulary, and complex categorization schemes.
In summary, while CoT prompting and explanation-driven learning highlight the value of LLM rationales, their unstable quality poses risks if applied directly. Our framework addresses this by treating rationales not as deterministic supervision but as trust indicators that dynamically adjust each instance’s training weight. This preserves interpretability and generalization benefits while ensuring stability and reliability in high-stakes domains such as patent classification.
@@ 2.3 Trustworthiness and Selective Learning
Ensuring the trustworthiness of LLM predictions has become a critical issue as LLMs are increasingly used in decision-sensitive contexts. A major line of work improves reliability at the inference time. Self-Consistency aggregates multiple reasoning paths for stable outputs (Wang et al., 2023), selective prediction methods allow abstention under uncertainty (Yoshikawa and Okazaki, 2023), and self-evaluation frameworks train models to assign confidence scores (Chen et al., 2023). Uncertainty-aware fine-tuning further mitigates hallucinations and stabilizes reasoning (Krishnan et al., 2024; Lin et al., 2023). While introducing effective safeguards, these methods remain largely confined to inference-level mechanisms.
Our work instead integrates trust indicators into the training process. Rather than filtering outputs post hoc, we use rationale quality to dynamically reweight training instances. Specifically, we extend Self-Consistency as a measure of rationale stability during learning, and complement it with Class Entailment Alignment and LLM Agreement Scoring to evaluate semantic coherence and external plausibility. This selective distillation strategy makes the training procedure itself reliability-aware.
Such integration is important in domains like patent classification, where low-quality rationales can propagate errors with legal and economic impact. By embedding trust indicators into distillation, our framework advances beyond inference-level reliability toward responsible training practices, ensuring model improvements are guided by explanations that are both stable and faithful.
@@ 3.2 Reasoning Quality Evaluation: Lightweight Trust Scoring
By averaging over the predicted label set $\hat{Y}$, CEA evaluates the degree to which the rationale is consistent with the label space. The Cooperative Patent Classification (CPC) system follows a hierarchical structure, where broader technical domains (e.g., G06F for computer technology) contain more specific subclasses (e.g., G06N for artificial intelligence). In CEA, we explicitly leverage this property by extracting the formal definitions of predicted CPC codes through an LLM and aligning them with the generated rationale. For multi-label cases, where a patent may belong to several subclasses simultaneously, CEA averages the alignment across all candidate CPC definitions, ensuring that the rationale remains coherent with the domain-specific taxonomy.
For instance, in patent classification, when a document is assigned to H04L (digital communication) but the generated rationale emphasizes aspects of “antenna structure” rather than “signal transmission protocols,” the semantic mismatch results in a low CEA score. This reflects misalignment between the rationale and the formal CPC definition, which our framework treats as an indicator of reduced trustworthiness.
@@ 3.3 Trust-Aware Training Control
$\mathcal{L}=\sum_{x\in\mathcal{D}}{1}\{\mathrm{CTS}(x)\geq\tau\}\cdot\mathrm{CTS}(x)\cdot\mathrm{Loss}\big(f_{\theta}(x),\,y\big).$ (6)
@@ 4 Experiments and Results
We evaluate the proposed framework by addressing four research questions that collectively assess its effectiveness, robustness, and interpretability in patent classification.
RQ1: How does SFD compare with baseline methods? RQ2: Does threshold-based filtering improve robustness compared to soft weighting? RQ3: What is the effect of removing each trust metric? RQ4: Are trust scores correlated with human judgments of rationale quality?
@@ 4.1 Dataset
In this study, we employ the USPTO-2M dataset (United States Patent and Trademark Office, ). USPTO-2M is a large-scale dataset released by the United States Patent and Trademark Office (USPTO), consisting of more than two million patent documents and thousands of CPC codes. This dataset has become one of the most widely used benchmarks for multi-label patent classification and has been consistently adopted as a reliable evaluation resource in numerous research efforts (Lee and Hsiang, 2019; Bekamiri et al., 2021; Zhou et al., 2021).
Patent documents differ substantially from general-domain texts due to their use of specialized terminology, complex sentence structures, and strong interdependencies among technical concepts. These characteristics introduce significant challenges for automated classification. To ensure both coverage and feasibility, we restrict the label space to the subclass level of the CPC hierarchy. This setting prevents over-fragmentation of labels while maintaining sufficient representativeness of the data, as widely adopted in prior patent classification studies (Lee and Hsiang, 2019; Haghighian Roudsari et al., 2022; Risch et al., 2020).
Therefore, USPTO-2M serves as an appropriate evaluation basis for this work. It not only reflects the inherent challenges of patent data such as large scale, complexity, and hierarchical structure, but also provides a robust foundation for fair comparison across different modeling approaches.
@@ 4.2 Experimental Settings
Qwen3-30B is used as the teacher model with Qwen3-0.6B as the student model in our distillation framework. For self-consistency, we generate three rationales per instance, balancing efficiency and stability. Additional rationales yield marginal gains with higher cost. Qwen3-32B serves as the Judge LLM for LAS computation, providing an independent evaluation of rationale–label plausibility. The trust coefficients $\lambda_{1},\lambda_{2},\lambda_{3}$ are fixed at $1/3$, treating each metric equally as in the definition of $\mathrm{CTS}(x)$. We optimize the student model using the Adam optimizer with a learning rate of 1e-5. All experiments are conducted on the USPTO-2M dataset, as described in Section 4.1.
@@
SFD 0.960 0.404
Table 1: Performance comparison of SFD against baseline methods on USPTO-2M. Results show improvements in both Micro-F1 and Macro-F1, indicating enhanced effectiveness and robustness.
We first examine whether the proposed SFD framework outperforms existing baseline approaches. Table 1 presents the experimental results on the USPTO-2M dataset.
The label-only distillation baseline shows the weakest performance with an F1-Micro of 0.821 and an F1-Macro of 0.401. Naive CoT distillation achieves a strong F1-Micro of 0.942, but its F1-Macro remains at 0.401, indicating limited improvement on minority classes. Self-consistency distillation and LLM-as-a-judge distillation record 0.932/0.354 and 0.893/0.355 respectively, showing partial gains but lacking balance across labels. Rationale-augmented distillation reaches 0.944/0.369, which is relatively stable yet still below the best result.
By contrast, SFD achieves the top scores with 0.960 and 0.404 (F1-Micro, F1-Macro), surpassing all baselines. The improvement in Macro-F1 highlights that the method delivers reliable reasoning signals for rare categories. These findings demonstrate that treating LLM-generated rationales as trust indicators and selectively incorporating them into training yields both performance gains and improved robustness in patent classification.
@@
Table 2: Results of threshold-based filtering on USPTO-2M, showing the trade-off between Micro-F1, Macro-F1, and Subset Accuracy across different $\tau$ values.
We compare two strategies: (i) soft weighting, where all samples contribute proportionally to their trust scores ($\tau=0$); and (ii) threshold-based filtering, which excludes samples with $\mathrm{CTS}(x)<\tau$. As shown in Table 2, mild filtering ($\tau>0.5$) yields limited improvement, while progressively stronger thresholds lead to more robust performance. At $\tau>0.8$, the model achieves the highest Macro-F1 (0.462), showing strong reliability for rare classes. At $\tau>0.9$, both Micro-F1 (0.960) and Subset Accuracy (0.930) reach their peak values, indicating that retaining only the most trustworthy rationales produces the most stable and reliable predictions overall. This demonstrates that high-threshold filtering effectively suppresses noisy reasoning while preserving informative signals. Consequently, we select $\tau>0.9$ as the optimal configuration for subsequent experiments.
@@
We conduct an ablation study on the USPTO-2M dataset to evaluate the relative contributions of each trust metric. Performance is measured using F1-Micro, F1-Macro, and Subset Accuracy. Following the results of RQ2, we fix the threshold at $\tau>0.9$ to ensure consistency across experiments.
[…]
These findings suggest that CEA provides the strongest domain-grounded signal, while SC and LAS complement CEA by enhancing stability and plausibility. Their incorporation is essential not only for maximizing predictive accuracy but also for ensuring that reasoning remains interpretable and trustworthy in patent classification tasks.
4.6 RQ4. Are trust scores correlated with human judgments of reasoning quality?
@@
Table 4: Pearson correlation ($\rho$) between trust scores and human judgments on 1,000 samples of USPTO-2M. $\Delta$ and Rel. $\Delta$ denote absolute and relative differences from $\mathrm{CTS}(x)$.
We sample 1,000 instances from USPTO-2M with varied trust scores. Three annotators rate each rationale on a 1–5 Likert scale for logical consistency, task alignment, and plausibility, with Krippendorff’s $\alpha=0.641$. Guidelines are in Appendix D.
Table 4 shows that the $\mathrm{CTS}(x)$ correlates strongest with human ratings, $\rho=0.685$. CEA is most critical, with its removal causing the largest drop at -30.66%, followed by LAS at -22.92%, and SC at -7.88%. $\mathrm{CTS}(x)$ enhances the interpretability in patent classification.
Overall, these findings demonstrate that all three metrics contribute meaningfully, with CEA playing the most critical role. $\mathrm{CTS}(x)$ provides a more reliable and interpretable signal of reasoning quality, aligning closely with human judgments in complex domains such as patent classification.

### v3

@@ Abstract
Large language models (LLMs) increasingly generate natural language rationales to enhance interpretability, but these often contain logical errors, label mismatches, and domain-specific misalignments. Directly using such rationales as supervision risks propagating noise and undermining training stability. To address this challenge, we introduce Self-Filtered Distillation, a framework tailored for patent classification that treats LLM-generated rationales as trust signals rather than ground-truth supervision. The framework employs selective distillation guided by three unsupervised trust metrics: (1) Self-Consistency, which measures the stability of LLM-generated rationales across multiple generations; (2) Class Entailment Alignment, which assesses semantic coherence with patent-specific class definitions; and (3) LLM Agreement Scoring, which validates rationale–label plausibility. These metrics are integrated into a unified trust score that primarily weights training samples while optionally filtering out extremely low-trust cases, enabling reasoning-aware supervision. Experiments on the USPTO-2M dataset show that our method consistently outperforms label-based learning and conventional distillation in accuracy, stability, and interpretability across diverse student architectures, establishing a reliable paradigm for leveraging reasoning-aware trust indicators in patent analytics.
@@ 1 Introduction
Patent classification is critical for organizing large-scale patent corpora into technical classification schemes, serving as the foundation for applications such as intellectual property management, technology scouting, and innovation analysis (jiang2024natural; yoo2025patentmind). Unlike general-domain texts, patents exhibit a multifaceted nature: highly specialized terminology, complex claim structures, and strategically ambiguous expressions, which pose significant challenges for automated classification. Our approach is specifically tailored to the patent domain, distinguishing it from general text classification techniques. Recent advances in large language models (LLMs) offer strong potential to address these challenges by processing massive patent datasets with improved accuracy and generating natural language rationales that enhance interpretability and provide a reasoning-based foundation for trust in classification outcomes (lee2019patentbert; bai2024patentgpt).
Figure 1: Overview of the Self-Filtered Distillation (SFD) strategy. The framework evaluates LLM-generated rationales using three trust indicators. Rationales with high trust scores are incorporated into the student model training via soft weighting proportional to their trust score, while those falling below the threshold $\tau$ are filtered out to maintain training stability.
@@
Beyond predictive accuracy, research in NLP has increasingly emphasized explanation-driven learning as a means of enhancing interpretability and trustworthiness. This line of work leverages human-annotated rationales as auxiliary supervision signals, demonstrating that explanations can improve both generalization and interpretability (camburu2018esnli; deyoung2020eraser). While promising, such approaches incur high annotation costs and are often limited to short texts and simple classification tasks. Consequently, they fall short when applied to the classification of multifaceted and lengthy patents, where explanation-driven supervision faces domain-specific challenges.
@@
With the advent of LLMs, research has shifted toward leveraging model-generated rationales to strengthen learning without requiring human annotation (wei2022chain; hsieh2023distilling). While this paradigm offers scalability and interpretability, LLM-generated rationales often contain logical inconsistencies, label mismatches, and semantic misalignment with domain-specific classification schemes (turpin2023language) that are unsuitable for patents. For instance, an LLM may justify categorizing a digital communication patent (H04L) based on antenna structures or misclassify an artificial intelligence patent (G06N) as general data processing (G06F). Such errors extend beyond performance degradation: in the patent domain, they can lead to misguided technology mapping, flawed market analysis, or even intellectual property disputes. Treating all rationales as direct supervisory signals risks propagating these errors and reducing training stability. Therefore, high-stakes applications of patent classification demand a framework that can quantitatively assess rationale quality and selectively incorporate only trustworthy rationales (lampinen2022can).
@@
Empirical evaluations on the USPTO-2M dataset demonstrate that SFD consistently surpasses traditional label-based learning and conventional distillation methods across diverse student architectures, achieving superior performance in predictive accuracy, training stability, and interpretability. These results indicate that incorporating unsupervised trust indicators as quality filters effectively enhances classification performance and establishes a statistically robust learning paradigm for leveraging LLM-generated rationales in complex, high-stakes domains.
@@
  • Trust-based distillation paradigm: treats LLM-generated rationales as trust indicators rather than ground-truth signals, reducing instability and error propagation while turning explanations into reliable learning resources.
@@
  • Patent-domain validation: Applied to patent classification, SFD outperforms label-based learning and conventional distillation in accuracy, stability, and interpretability, demonstrating value for trustworthy patent analytics.
@@ 2.1 Patent Classification
Patent classification is a central task in intellectual property management, supporting prior art search, and legal decision-making. Early studies used TF–IDF features with classifiers such as SVMs or logistic regression, with limited performance due to domain-specific terminology and high-dimensional feature spaces (fall2003automated).
Deep learning brings about significant improvements: CNN- and RNN-based models capture semantic and structural patterns from claims and abstracts, outperforming methods (li2018deeppatent). Still, the challenges of long documents, technical vocabulary, and concept interdependencies make patent classification harder than general text classification. Transformer-based models open a new paradigm. Domain-adapted encoders such as PatentBERT (lee2019patentbert) and PatentSBERTa (bekamiri2021patentsberta) outperform BERTs, while generative approaches like PatentGPT show the potential of large-scale pretraining (bai2024patentgpt). Leveraging the hierarchical structure of the Cooperative Patent Classification (CPC) also helps address the complexity of multi-label classification (zhou2021hierarchical). Large-scale datasets like USPTO-2M, with millions of patents and thousands of CPC classes, further accelerate progress but also introduce issues of computational cost and severe label imbalance (risch2020patentmatch).
@@ 2.2 Explanation-driven Learning
Chain-of-Thought (CoT) prompting enhances LLM performance and interpretability in complex reasoning tasks by generating intermediate rationales (wei2022chain; kojima2022zero; wang2023selfconsistency). This has inspired distillation methods where student models learn from teacher reasoning traces (hsieh2023distilling). While approaches like KeyPoint-CoT improve efficiency (feng2024keypointcot), most assume uniform rationale quality. This risks propagating errors, a critical issue in patent classification where misaligned rationales can distort label assignments. Explanation-driven learning utilizes rationales as auxiliary supervision, with human-annotated datasets demonstrating improved generalization and robustness (camburu2018esnli; deyoung2020eraser; lampinen2022can; wiegreffe2021measuring). Although metrics like ROSCOE exist to assess faithfulness (golawn2022roscoe; hase2020evaluating), LLM rationales often suffer from incompleteness and label inconsistencies (turpin2023language; ye2022unreliability; zhu2025rationales). These flaws are amplified in patent classification due to its complex technical vocabulary and categorization schemes.
While CoT and explanation-driven learning highlight the utility of rationales, their instability poses risks if applied directly. Our framework addresses this by repurposing evaluation concepts, such as consistency and alignment, as trust indicators to dynamically adjust training weights. This approach preserves interpretability benefits while ensuring stability and reliability in high-stakes domains like patent classification.
@@ 2.3 Trustworthiness and Selective Learning
Ensuring LLM trustworthiness is critical in decision-sensitive contexts. Existing work largely focuses on inference-time reliability through methods like Self-Consistency (wang2023selfconsistency) and semantic uncertainty quantification (kuhn2023semantic). Other approaches employ selective prediction or self-evaluation for confidence scoring (yoshikawa2023selective; chen2023selfeval). While uncertainty-aware fine-tuning helps mitigate hallucinations (krishnan2024enhancing; lin2023uncertainty; damani2025beyond), these safeguards remain confined to inference or require specific tuning objectives. Research on self-improvement explores bootstrapping via synthetic data, often leveraging Reinforcement Learning (RL) (zelikman2022star; huang2023selfimprove). However, these powerful methods typically rely on ground-truth verification or computationally expensive reward models to filter outputs.
Our work integrates trust indicators into the training process as a more efficient alternative. Instead of post-hoc filtering or RL loops, we dynamically reweight training instances based on rationale quality using Self-Consistency, Class Entailment Alignment, and LLM Agreement Scoring. This selective distillation strategy enables reliability-aware training without expensive annotations. In high-stakes domains like patent classification, this prevents error propagation from low-quality rationales, advancing beyond inference-level reliability toward responsible, stability-guided training.
3 Methodology
@@ 3.2 Reasoning Quality Evaluation: Lightweight Trust Scoring
By averaging over the predicted label set $\hat{Y}$, CEA evaluates the rationale’s consistency with the label space. The Cooperative Patent Classification (CPC) system follows a hierarchical structure, where broader technical domains (e.g., G06F for computer technology) contain specific subclasses (e.g., G06N for artificial intelligence). CEA leverages this by extracting definitions of predicted CPC codes via an LLM and aligning them with the rationale. For multi-label cases, CEA averages alignment across all candidate definitions, ensuring coherence with the domain taxonomy. For instance, when a document is assigned to H04L (digital communication) but the rationale emphasizes “antenna structure” rather than “signal transmission protocols,” the semantic mismatch results in a low CEA score, indicating reduced trustworthiness.
@@ 3.3 Trust-Aware Training Control
$\mathcal{L}=\sum_{x\in\mathcal{D}}\mathbf{1}\{\mathrm{CTS}(x)\geq\tau\}\cdot\mathrm{CTS}(x)\cdot\mathrm{Loss}\big(f_{\theta}(x),y\big)$
@@ 4 Experiments and Results
We evaluate the proposed framework by addressing five research questions that collectively assess its effectiveness, robustness, and interpretability in patent classification. RQ1: How does SFD compare with baseline methods? RQ2: Does threshold-based filtering improve robustness compared to soft weighting? RQ3: What is the effect of removing each trust metric? RQ4 : Is the proposed framework model-agnostic across different student architectures? RQ5: Are trust scores correlated with human judgments
@@ 4.1 Dataset
We employ the USPTO-2M dataset (uspto2m), a predominant benchmark in multi-label patent classification comprising over two million documents. This extensive scale encompasses diverse technical descriptions across various fields, effectively mitigating selection bias and ensuring statistically robust evaluation. Patent texts exhibit unique challenges, including specialized terminology, complex structures, and concept interdependencies, which distinguish them from general-domain texts. To balance coverage with computational feasibility, we restrict the label space to the subclass level of the Cooperative Patent Classification (CPC) hierarchy. This configuration prevents label over-fragmentation while preserving sufficient data representativeness, a standard practice in prior studies (lee2019patentbert; patentnet; Risch2020HierarchicalSeqGen). USPTO-2M serves as a robust and rigorous foundation for evaluating model performance against these inherent domain challenges.
@@ 4.2 Experimental Settings
We employ Qwen3-30B as the teacher and validate our framework across four diverse student models: Qwen3-0.6B, Phi-4-Mini, Llama-3.2-3B, and Falcon3-3B. Rationales are generated via Nucleus sampling ($T=0.6,p=0.9$), with results averaged over five independent runs to ensure reliability. For trust evaluation, we sample three rationales per instance and set equal coefficients ($\lambda=1/3$). We utilize Qwen3-32B as the judge for its strong legal reasoning capabilities and all-MiniLM-L6-v2 for embeddings. Training uses the Adam optimizer with a learning rate of 1e-5. All code and checkpoints will be released upon acceptance.
@@
Our Framework (SFD)
SFD w/ Qwen3-0.6B 0.960 0.404
SFD w/ Phi-4-Mini 0.969 0.503
SFD w/ Llama-3.2-3B 0.988 0.515
SFD w/ Falcon3-3B 0.988 0.556
Table 1: Performance comparison on USPTO-2M. Our SFD framework consistently outperforms strong baselines, achieving up to a 38.7% relative improvement.
Table 1 compares the performance of SFD against various distillation baselines on the USPTO-2M dataset. A critical observation is that SFD consistently outperforms the strongest baselines across all tested configurations. To rigorously validate the framework’s robustness, we conducted five independent runs for all student architectures. Every model demonstrated stable performance gains with minimal variance; for instance, SFD achieved $0.960\pm 0.006$ Micro-F1 with Qwen3-0.6B and $0.988\pm 0.003$ with Llama-3.2-3B. These results confirm that our trust-aware mechanism effectively stabilizes training regardless of the underlying backbone. Furthermore, SFD achieved a remarkable breakthrough with the Falcon3-3B backbone, recording a Micro-F1 of 0.988 and a Macro-F1 of 0.556, which represents a relative improvement of up to 38.7% over the strongest baseline. These results collectively demonstrate that SFD not only enhances predictive accuracy but also ensures high reliability in complex patent classification tasks.
@@
Table 2: Effect of threshold-based filtering using Qwen3-0.6B on USPTO-2M. Progressively stricter filtering improves robustness, with $\tau>0.9$ achieving a peak Subset Accuracy of 0.930.
To evaluate the impact of filtering, we use Qwen3-0.6B as the student model. We compare soft weighting, where all samples contribute proportionally, against threshold-based filtering which excludes samples with low trust scores. As shown in Table 2, mild filtering yields limited improvement, whereas progressively stronger thresholds lead to substantial performance gains. Notably, setting the threshold at $\tau>0.9$ achieves peak performance for the Qwen3-0.6B model, recording 0.960 in Micro-F1 and 0.404 in Macro-F1, along with a significant boost in Subset Accuracy to 0.930. This indicates that retaining only the most trustworthy rationales effectively suppresses noisy reasoning while preserving informative signals. Consequently, we select $\tau>0.9$ as the optimal configuration for subsequent experiments.
@@
We conducted an ablation study on the USPTO-2M dataset using Qwen3-0.6B to rigorously evaluate the contribution of each trust metric, with the threshold fixed at $\tau>0.9$. As shown in Table 3, the integrated trust score ($\mathrm{CTS}(x)$) yields the most robust performance, achieving 0.960 in Micro-F1 and 0.404 in Macro-F1. A hierarchical analysis of the metrics reveals distinct roles for each component. Removing CEA results in the most severe degradation, causing Subset Accuracy to plummet to 0.695. This confirms that enforcing semantic alignment with CPC definitions is the foundational requirement for correct classification. The exclusion of LAS leads to a significant drop in overall precision, indicating its critical function in filtering out plausible but hallucinated rationales. Interestingly, removing SC has a marginal impact on Micro-F1 (0.957) but notably affects Macro-F1 (0.359). This suggests that while self-consistency is less critical for clear-cut cases, it plays a vital role in stabilizing predictions for complex or minority classes. Therefore, while CEA provides the necessary domain grounding, the complementary integration of LAS and SC is essential for maximizing both reliability and generalization.
4.6 RQ4. Is the proposed framework model-agnostic across different student architectures?
Student Model Micro-F1 Macro-F1 Subset Acc.
Qwen3-0.6B 0.9600 0.4040 0.9330
Phi-4-Mini 0.9685 0.5025 0.9453
Llama-3.2-3B 0.9880 0.5153 0.9765
Falcon3-3B 0.9881 0.5556 0.9766
Table 4: Performance comparison across different student architectures using the SFD framework.
To verify whether the effectiveness of SFD is independent of specific model structures, we evaluated the framework across a diverse set of student architectures, including Qwen3, Phi-4, Llama-3.2, and Falcon3. As shown in Table 4, SFD consistently delivers robust performance across all tested backbones, confirming that the benefits of trust-aware distillation are not limited to a single model type. All models achieved high accuracy, with Micro-F1 scores exceeding 0.960 across the board. Notably, the Falcon3-3B model achieved the highest performance, with a Micro-F1 0.988 and a Macro-F1 0.556. These results demonstrate that our framework effectively transfers reasoning capabilities to various architectures, proving its generalizability and model-agnostic nature.
4.7 RQ5. Are trust scores correlated with human judgments of reasoning quality?
@@
Table 5: Pearson correlation ($\rho$) between trust scores and human judgments on 1,000 samples of USPTO-2M. $\Delta$ and Rel. $\Delta$ denote absolute and relative differences.
We sampled 1,000 instances from USPTO-2M with varied trust scores. Three annotators rated each rationale on a 1–5 Likert scale for logical consistency, task alignment, and plausibility, with Krippendorff’s $\alpha=0.641$. Guidelines are in Appendix D. Table 5 shows that the $\mathrm{CTS}(x)$ correlates strongest with human ratings, $\rho=0.685$. CEA is most critical, with its removal causing the largest drop at -30.66%, followed by LAS at -22.92%, and SC at -7.88%. $\mathrm{CTS}(x)$ enhances interpretability in patent classification. Overall, these findings demonstrate that all three metrics contribute meaningfully, with CEA playing the most critical role. Crucially, the dominance of CEA, which anchors evaluation to objective ontology definitions, combined with the high correlation with human experts ($\rho=0.685$), confirms that our trust scoring is robust against potential model-family preference bias. Consequently, $\mathrm{CTS}(x)$ provides a reliable and interpretable signal of reasoning quality, aligning closely with human judgments in complex domains such as patent classification.